OpenStudy (anonymous):

HELP PLEASE - I DONT UNDERSTAND The length of a rectangular suitcase is 7 inches longer than the width. The area of the suitcase is 144 squared inches. Find the dimensions of the suitcase. I KNOW LENGTH=WIDTH+7

OpenStudy (anonymous):

So if Area = Length x width then you substitute (width+7) for Length in the equation what would the equation look like?

OpenStudy (anonymous):

IM SORRY YOUR NOT HELPING ; YOUR MAKING ME MORE CONFUSED THEN I ALREADY AM

OpenStudy (anonymous):

IF 10 = 2x +2y and you know that y = 3x..how would you solve?

OpenStudy (anonymous):

IM SORRY YOUR NO HELP IM SORRY .

OpenStudy (tanya123):

@iPwnBunnies best helper

OpenStudy (anonymous):

THANK YOU @tanya123

OpenStudy (tanya123):

THANKS FOR MAKING ME A HUMAN CALCULATOR !!!

geerky42 (geerky42):

PLEASE DO NOT TYPE LIKE THIS! IT MAKES YOU LOOK LIKE YELLING

geerky42 (geerky42):

:(

OpenStudy (ipwnbunnies):

They just ask for the area of the suitcase. Area = length*width The length is 7 inches longer than the width. Let w be the width. Length will be (w + 7) A = w(w+7)

OpenStudy (anonymous):

sorry i aint yelling @geerky42

OpenStudy (ipwnbunnies):

144 = w(w+7), solve for w. The width of the suitcase

OpenStudy (anonymous):

would it be 16 and 9 ? @iPwnBunnies

geerky42 (geerky42):

ya 16in by 9in good work! :)

OpenStudy (anonymous):

thank you